On June 5, 2017, Javier Sanchez purchased and placed in service a new

7-year class asset costing $560,000 for use in his landscaping business,

which he operates as a single member LLC (Sanchez Landscaping LLC). During

2017, his business generated a net income of $945,780 before any§ 179 immediate

expense election.

a. Determine the maximum deductions that Javier can claim w ith respect to this

asset in 2017 and 2018.
Answer #1

Year 2017

The amount of the deduction is the total cost of the Asset upto $500,000 for each individual item.

Total cost                                                    560,000

Sec 179 Deduction                                    -500,000

Bonus Depreciation Deduction                    -30,000

(50% CAP in 2017)

Normal 1st year Depreciation (60000/7)=       -8570

Total Deduction in 2017 : (500,000+30000+12000) = 538570

Year 2018

60000/7 =   8570